OpenStudy (anonymous):

If the perimeter of a rectangle is 152m and the length is 7m more than twice the width how do I find the length and width?

OpenStudy (anonymous):

P =(L+W)*2 L=7+2W 152= (7+2W)+W))*2 152/2= (7+2W)+W 76 = 7+ 3W 76-7 = 3W 69/3 = W 23= W L = 7+(2*23) L =53

OpenStudy (anonymous):

Let L be the length and W be the width. Well the perimeter is 2L + 2W, so 2L + 2W = 152 And the length is given to be 7 more than 2 times the width so: 2W +7 = L. Therefore 2W = L - 7 Putting both those equations together you have: \[2L + 2W = 152\]\[\implies 2L + (L - 7) = 152\]\[\implies 3L = 152 + 7\]\[\implies 3L = 159 \]\[\implies L = 53\] And since \[2W = L - 7\]\[\implies 2W = 53 - 7 = 46\]\[\implies W = 46/2 = 23\]

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

u r w come